The stopping distance D of a car after the brakes are applied varies directly as the square of the speed R. If a car traveling 50 mph can stop in 140 ft, how many feet will it take for the same car to stop traveling 90 mph?
Please help with this word problem. Thank you.

Answer by checkley77(12844)

50^2/140=90^2/x
50^2x=8,100*140
2,500x=1,134,000
x=1,134,000/2,500
x=453.6 feet.